# Learning Equivariant Maps with Variational Quantum Circuits

### v1

@@ II Symmetry Assisted Quantum Machine Learning
This lemma gives us a technique for constructing VQCs which respect the symmetry of a problem. First find a representation $W$ induced by the embedding $\mathcal{E}$. Construct a $G$-invariant gate set by twirling self-adjoint operators and using them as the infinitesimal generators of the gate set. Finally, choose an observable which is $G$-invariant. By Proposition 1, the function estimate produced by the model will now be $G$-invariant. This approach is demonstrated in the next section.
III Classifying Lines in Images
Let us demonstrate how to construct VQCs which respect the symmetry of a problem. Consider the problem of classifying vertical and horizontal lines in $2\times 2$ grayscale images (see Figure 1).
An image of a vertical line is still an image of a vertical line when it is rotated by 180 degrees. Similarly, the orientation of a horizontal line is preserved by this rotation. In fact, we also preserve the labeling when the image is flipped about a vertical line through its midpoint. Each of these operations has order two since they undo themselves; thus, this image classification problem has symmetry group $G=C_{2}\times C_{2}$, where $C_{2}$ is the cyclic group of order two.
@@ III Classifying Lines in Images
The simplest nontrivial encoding for this problem is the amplitude encoding, wherein a data point $(a,b,c,d)^{T}$ is mapped to the quantum state with the same coefficients (up to normalization) in the computational basis:
@@
With this encoding, the induced representation is the same as the representation $V$, so that $G$-invariant operators need only commute with $I\otimes X$ and $X\otimes X$. Equivalently, we could take the vertical and horizontal flips to be the generators of $C_{2}\times C_{2}$, so that a $G$-invariant operator need only commute with $I\otimes X$ and $X\otimes I$. Clearly the generators themselves satisfy this property since $(I\otimes X)(X\otimes I)=X\otimes X=(X\otimes I)(I\otimes X)$. Thus, the rotations by an arbitrary angle about these axes are also $G$-invariant by Lemma 1, and so we can use them to parametrize our circuit. These rotations correspond to $I\otimes R_{X}(\theta)$ and $R_{X}(\theta)\otimes I$. In fact, by invoking our commutation relations, it is easily shown that the only two-qubit $G$-invariant operators have the form $aI\otimes I+bI\otimes X+cX\otimes I+dX\otimes X$ for some constants $a,b,c,d$. Since each of the terms commutes with the rest, the exponentiation splits into a product of four exponentials by the Baker-Campbell-Hausdorff formula. That is, the exponential $\exp(i\theta(aI\otimes I+bI\otimes X+cX\otimes I+dX\otimes X))$ is equivalently given by
$e^{i\theta aI\otimes I}e^{i\theta bI\otimes X}e^{i\theta cX\otimes I}e^{i\theta dX\otimes X}.$ (28)
@@
Next, we will adjust this encoding slightly by introducing a readout qubit. The encoding can now be thought of as the tensor product of the amplitude encoding with an identity mapping on the readout qubit, which is initialized to |$\ket{0}$. The induced representation $W$ is now defined by
$\left(\mathcal{E}_{amp}\otimes I\right)\circ V(g)=W(g)\circ\left(\mathcal{E}_{amp}\otimes I\right),$ (29)
so that $V$ and $W$ are again identical, although this time they act on three qubits. We must find self-adjoint operations that commute with $X\otimes I\otimes I$ and $I\otimes X\otimes I$. Clearly $I\otimes I\otimes H$ satisfies this property for any single qubit self-adjoint operator $H$. In particular, we see that we may make our measurement with respect to $Z$ on the readout qubit at the end of the circuit while still satisfying the hypothesis of Proposition 1. The $X$ gate commutes with these operations, no matter which qubit it acts on, so we are free to use $R_{x}(\theta)$ gates in the construction of the variational circuit. The self-adjoint operation $I\otimes I\otimes Z+I\otimes X\otimes X$ also commutes with both $X\otimes I\otimes I$ and $I\otimes X\otimes I$. Moreover, $[Z,X]\neq 0$, so that the exponentiation of $I\otimes I\otimes Z+I\otimes X\otimes X$ is not a product of local operations and can produce entanglement between a data qubit and the readout qubit. A similar conclusion can be made for the operation $I\otimes I\otimes Z+X\otimes I\otimes X$. The circuit is therefore made symmetric by measuring the readout qubit with respect to $Z$ (or any other single qubit unitary) and building the variational circuit from the gate set $\{R_{X}(\theta),e^{i\theta\left(I\otimes I\otimes Z+I\otimes X\otimes X\right)},e^{i\theta\left(I\otimes I\otimes Z+X\otimes I\otimes X\right)}\}$.
An example circuit for solving the horizontal/vertical line classification problem with a built in symmetry bias is shown in Figure 2. This circuit has a single trainable parameter and solves the problem with 100% accuracy after training.
@@
IV Learning Equivariant Maps
Given the representation $V:G\to GL(\mathcal{R})$ describing the symmetry respected by the data space and another representation $W:G\to\mathcal{U}(\mathcal{H})$, we can learn an equivariant map between them, assuming it exists. In the context of the symmetry assisted VQC outlined above, this equivariant map appears as the data embedding, so that we are learning a data embedding which properly respects the symmetry of the data. We begin by encoding the data with a parameterized unitary $\mathcal{E}_{\theta}(x)$ and follow this with a unitary which is $G$-invariant with respect to $W$. The expected value of the circuit will then be $G$-invariant with respect to $V$ whenever the parameterized unitary is an equivariant map between $V$ and $W$ and the observable is $G$-invariant with respect to the $W$ representation. To train the equivariance, we sample from both the data space and the group and compute the loss function as $L_{g}(x)=(\mathcal{E}_{\theta}(V(g)x)-\mathcal{E}_{\theta}(x))^{2}$. We then employ the Simultaneous Perturbation Stochastic Approximation (SPSA) [30] method to perform gradient descent. Thus, learning an equivariant map amounts to performing a regression minimizing $L_{g}(x)$ for any pair $(g,x)\in G\times\mathcal{R}$. It should be noted that the value of $L_{g}(x)$ can be interpreted as a measure of how equivariant the embedding $\mathcal{E}_{\theta}$ is; that is, how well it transfers the notion of symmetry from one representation to another. In most cases, $L_{g}(x)$ is not actually zero, and so we are really learning maps which are in a sense almost equivariant. We will now illustrate this construction for a variety of example symmetries.
IV.1 Example: $C_{2}$ Symmetry
@@
Consider again the problem of classifying vertical and horizontal lines in a $2\times 2$ image and recall that the classification of a line is invariant under a flip across the vertical axis. Since this flip is its own inverse operation, the group it generates is the cyclic group of order two. Thus, we have a representation $V$ of $C_{2}$ in the dataspace. On the other hand, the operation $\textnormal{SWAP}_{01}\textnormal{SWAP}_{23}$ generates a representation $W$ of $C_{2}$ in the Hilbert space setting and we can attempt to find an equivariant map between these representations by introducing a parameterized embedding $\mathcal{E}_{\theta}(x)=R_{Y}(\theta_{0}+x_{0})\otimes R_{Y}(\theta_{1}+x_{1})\otimes R_{Y}(\theta_{2}+x_{2})\otimes R_{Y}(\theta_{3}+x_{3})$, where $\theta=(\theta_{0},\theta_{1},\theta_{2},\theta_{3})$ and $x=(x_{0},x_{1},x_{2},x_{3})$. Note that the architecture of the embedding is a free parameter; we have a choice of what gates to allow in its construction. Here we choose a tensor product of rotation gates to easily showcase our procedure. Let $g$ be the generator of $C_{2}$ and observe that $\mathcal{E}_{\theta}\circ V(g)(x)$ is given by
@@
We see that the periodicity of the rotation operators allows for several choices of the $\theta_{i}$ which produce equivariance. For example, when $\theta_{0}=\theta_{1}+2\pi n$ and $\theta_{2}=\theta_{3}+2\pi m$ for some $m,n\in\mathbb{Z}$, it follows that $\mathcal{E}_{\theta}$ is an equivariant map between the representations. Alternatively, if each of the pairs $\theta_{0},\theta_{1}$ and $\theta_{2},\theta_{3}$ differs by an odd multiple of $\pi$, the resulting signs cancel and we again achieve equivariance. In our simulation we measure with respect to the observable $X\otimes X\otimes Z\otimes Z$, which commutes with the operation $\textnormal{SWAP}_{01}\textnormal{SWAP}_{23}$ and is therefore $G$-invariant, and find that the model learns the parameters $\theta_{0}\approx 3.07\approx\pi$, $\theta_{1}\approx 6.18\approx 2\pi$, $\theta_{2}\approx 4.71\approx 3\pi/2$, and $\theta_{3}\approx 1.65\approx\pi/2$, which aligns with what we expect. The circuit is shown in Figure 3, the parameter curves are shown in Figure 4, and a plot of the average validation loss is shown in Figure 5.
IV.2 Example: $C_{2}\times C_{2}$ Symmetry
Let us add an independent symmetry to the mix. The flip $f_{H}$ about the horizontal axis in the $2\times 2$ line classification problem is independent from the vertical flip $f_{V}$, and we will now endeavour to take both into account. The group generated by these independent flips is a direct product of cyclic groups $C_{2}\times C_{2}$. We will make use of the same embedding as for the $C_{2}$ symmetry: $\mathcal{E}_{\theta}(x)=R_{Y}(\theta_{0}+x_{0})\otimes R_{Y}(\theta_{1}+x_{1})\otimes R_{Y}(\theta_{2}+x_{2})\otimes R_{Y}(\theta_{3}+x_{3})$, but this time we expect the parameters to train differently. Indeed, the representation of $C_{2}\times C_{2}$ acting on our Hilbert space will be given by $W(f_{V})=\textnormal{SWAP}_{01}\textnormal{SWAP}_{23}$ and $W(f_{H})=\textnormal{SWAP}_{02}\textnormal{SWAP}_{13}$, so that the unitary applied after the embedding must commute with these two operators. Of course, any choice of a single qubit gate applied simultaneously to all qubits will do the trick. Thus, we take our $C_{2}\times C_{2}$-invariant unitary to be $U=Y^{\otimes 4}$ and we take our observable to be $Z^{\otimes 4}$.
@@ IV.2 Example: $C_{2}\times C_{2}$ Symmetry
[b]
Figure 7: Parameter curves for $C_{2}\times C_{2}$ symmetry (a) and average loss for $C_{2}\times C_{2}$ symmetry (b).
Using the same reasoning as for the $C_{2}$ case, we expect to learn parameters which are equivalent up to some possible phase differences which cancel overall. This is indeed what we find, as the model learns the parameters $\theta_{0}\approx 3.24\approx\pi$, $\theta_{1}\approx 6.22\approx 2\pi$, $\theta_{2}\approx 4.89\approx 3\pi/2$, and $\theta_{3}\approx 4.97\approx 3\pi/2$. The latter two parameters are roughly equivalent, while the former two parameters are about $\pi/2$ from the latter parameter values in opposite directions, thereby canceling. See Figure 6 for the circuit, Figure 7(a) for the parameter curves, and Figure 7(b) for a plot of the average validation loss.
IV.3 Example: $D_{4}$ Symmetry
Consider a $3\times 3$ checkerboard in the plane–a generalization of the common XOR classification problem. This distribution demonstrates symmetry with respect to the dihedral group $D_{4}$ of order 8. In Figure 8, we plot several sample points which are separated into classes in alternating squares, forming just such a grid. Observe that a rotation $r$ by 90 degrees leaves this classification invariant, as does a flip $f$ about the diagonal line $y=x$. These symmetries together generate $D_{4}$, and so we take this to be our first representation.
@@ IV.3 Example: $D_{4}$ Symmetry
For the second representation, we will take the standard representation on four qubits. This representation is obtained by first identifying $D_{4}$ with the subgroup of the symmetric group $S_{4}$ generated by the permutations (1 2 3 4)1234(1\ 2\ 3\ 4)( 1 2 3 4 ) and (0 1)(2 3)0123(0\ 1)(2\ 3)( 0 1 ) ( 2 3 ) and then identifying these permutations with the corresponding permutation operators which act on four qubits. Thus, for the second representation we take $W(r)=\textnormal{SWAP}_{23}\textnormal{SWAP}_{12}\textnormal{SWAP}_{01}$ and $W(f)=\textnormal{SWAP}_{01}\textnormal{SWAP}_{23}$. Some operators which commute with these generators and are therefore $D_{4}$-invariant include a tensor product of single qubit operators $U^{\otimes 4}$ as well as the permutation $\textnormal{SWAP}_{02}\textnormal{SWAP}_{13}$. Let us therefore apply these operations with $U=Y$ and measure with respect to $Z^{\otimes 4}$. For the architecture of the embedding, we apply $R_{Y}(\theta_{i}+x_{i\!\!\mod 2})R_{Z}(\theta_{i+4}+x_{(i+1)\!\!\mod 2})$ to the $i$-th qubit. See Figure 9 for the circuit.
@@
The parameter curves and validation loss are shown in Figures 10 and 11, respectively. The resulting embedding after training is nearly equivariant. For example, taking the data point $(0.40,0.54)$, its rotation $\approx(2.46,0.40)$, and its reflection $\approx(0.54,0.40)$, and running the circuit with the trained parameters produces the nearly identical expected values $-0.05429996$, $-0.05429986$, and $-0.05430002$, respectively.
IV.4 Example: $S_{6}$ Symmetry
Consider the symmetric group $S_{6}$ on 6 letters. A problem which possesses this symmetry is the classification of graphs with 6 nodes as being either connected or disconnected. A graph is a collection of vertices $V$ and edges $E\subset V\times V$. When $(v,w)\in E$ if and only if $(w,v)\in E$ holds for every $v,w\in V$, we say that the graph is undirected. Otherwise, the graph is called directed or a digraph. We say that a given graph is connected if it has the property that for any vertices $v,w$, there is a sequence of edges $(v_{i},w_{i})$ with $i=0,1,\ldots,n$ such that $v=v_{0}$, $w_{n}=w$, and $w_{i}=v_{i+1}$ for all $i=0,\ldots,n-1$. Intuitively, the graph is connected if there is a path connecting any two vertices. Notice that the labeling of the vertices is arbitrary. We may choose to relabel the vertices with a permutation $\varphi:V\to V$ so long as vertices which are connected by an edge before the permutation remain connected after the permutation (equivalently, the edges are relabeled in precisely the same way). That is, we require that $\varphi$ is a bijection with the property that $(\varphi(v),\varphi(w))$ is an edge if and only if $(v,w)$ is an edge. Such a map is called a graph isomorphism and it can be shown that it preserves the property of connectedness.
@@ IV.4 Example: $S_{6}$ Symmetry
Figure 13: Average loss for $S_{6}$ symmetry.
Since permutations of the vertices are equivalent to graph isomorphisms, we take the collection of such maps as our representation of $S_{6}$ in the data space. In the Hilbert space, we take the standard representation of $S_{6}$ given by permuting tensor products according to the given permutation in $S_{6}$. Since this group is generated by the elements (1 2)12(1\ 2)( 1 2 ) and (1 2 3 4 5 6)123456(1\ 2\ 3\ 4\ 5\ 6)( 1 2 3 4 5 6 ), we see that the representation in Hilbert space is given by mapping (1 2)12(1\ 2)( 1 2 ) to $\textnormal{SWAP}_{01}$ and mapping (1 2 3 4 5 6)123456(1\ 2\ 3\ 4\ 5\ 6)( 1 2 3 4 5 6 ) to $\textnormal{SWAP}_{45}\textnormal{SWAP}_{34}\textnormal{SWAP}_{23}\textnormal{SWAP}_{12}\textnormal{SWAP}_{01}$. Certainly a tensor power $U^{\otimes 6}$ of any single qubit operator $U$ commutes with these generators and is therefore $S_{6}$-invariant. For this reason, we take $Z^{\otimes 6}$ as our observable and apply the operation $U=Y^{\otimes 6}$ after the encoding, which we give the following simple architecture: if $(i,j)$ is an edge in the graph with $i\neq j$, apply a controlled rotation with angle $\theta_{i}$ controlled off of qubit $i$ and targeting qubit $j$, and if $i=j$, then simply apply $R_{Y}(\theta_{i})$ to qubit $i$. The circuit is shown in Figure 14 for a particular graph defined by the notation {0: [3, 0, 5, 4], 1: [1, 5, 3, 0, 2], 2: [5, 3, 2, 4], 3: [5], 4: [4, 2], 5: [2, 5, 1, 3]}, which indicates, for example, that there is an edge between vertex 0 and vertices 3, 0, 5, and 4.
After training, the parameters converged to the same value, as seen in Figure 12, and this is what we expect since any relabeling of the vertices (maintaining adjacency) gives rise to an equivalent graph. The validation loss curve is plotted in Figure 13.

### v2

@@ II.1 Example: Classifying Lines in Images
An image of a vertical line is still an image of a vertical line when it is rotated by $180\degree$. Similarly, the orientation of a horizontal line is preserved by this rotation. In fact, we also preserve the same labeling when the image is flipped about a vertical line through its midpoint. Each of these operations has order two since they undo themselves; thus, this image classification problem has symmetry group $G=C_{2}\times C_{2}$, where $C_{2}$ is the cyclic group of order two.
@@
The simplest nontrivial embedding for this problem is the amplitude embedding, wherein a data point $(a,b,c,d)^{T}$ is mapped to the quantum state with the same coefficients (up to normalization) in the computational basis:
@@
With this embedding, the induced representation is the same as the representation $V$, so that $G$-invariant operators need only commute with $I\otimes X$ and $X\otimes X$. Equivalently, we could take the vertical and horizontal flips to be the generators of $C_{2}\times C_{2}$, so that a $G$-invariant operator need only commute with $I\otimes X$ and $X\otimes I$. Clearly the generators themselves satisfy this property since $(I\otimes X)(X\otimes I)=X\otimes X=(X\otimes I)(I\otimes X)$. Thus, the rotations by an arbitrary angle about these axes are also $G$-invariant by Lemma 1, and so we can use them to parametrize our circuit. These rotations correspond to $I\otimes R_{X}(\theta)$ and $R_{X}(\theta)\otimes I$. In fact, by invoking our commutation relations, it is easily shown that the only two-qubit $G$-invariant operators have the form $aI\otimes I+bI\otimes X+cX\otimes I+dX\otimes X$ for some constants $a,b,c,d$. Since each of the terms commutes with the rest, the exponentiation splits into a product of four exponentials by the Baker-Campbell-Hausdorff formula. That is, the exponential $\exp(i\phi(aI\otimes I+bI\otimes X+cX\otimes I+dX\otimes X))$ is equivalently given by
$e^{i\phi aI\otimes I}e^{i\phi bI\otimes X}e^{i\phi cX\otimes I}e^{i\phi dX\otimes X}.$ (26)
@@
Next, we will adjust this embedding slightly by introducing a readout qubit. The embedding can now be thought of as the tensor product of the amplitude embedding with an identity mapping on the readout qubit, which is initialized to |$\ket{0}$. The induced representation $W$ is now defined by
$\left(\mathcal{E}_{amp}\otimes I\right)(V(g)x)\ket{\vec{0}}=W(g)\left(\mathcal{E}_{amp}\otimes I\right)(x)\ket{\vec{0}},$ (27)
so that $V$ and $W$ are again identical, although this time they act on three qubits. We must find self-adjoint operations that commute with $X\otimes I\otimes I$ and $I\otimes X\otimes I$. Clearly $I\otimes I\otimes H$ satisfies this property for any single qubit self-adjoint operator $H$. In particular, we see that we may make our measurement with respect to $Z$ on the readout qubit at the end of the circuit while still satisfying the hypothesis of Proposition 1. The $X$ gate commutes with these operations, no matter which qubit it acts on, so we are free to use $R_{x}(\theta)$ gates in the construction of the variational circuit. The self-adjoint operation $I\otimes I\otimes Z+I\otimes X\otimes X$ also commutes with both $X\otimes I\otimes I$ and $I\otimes X\otimes I$. Moreover, $[Z,X]\neq 0$, so that the exponentiation of $I\otimes I\otimes Z+I\otimes X\otimes X$ is not a product of local operations and can produce entanglement between a data qubit and the readout qubit. A similar conclusion can be made for the operation $I\otimes I\otimes Z+X\otimes I\otimes X$. The circuit is therefore made symmetric by measuring the readout qubit with respect to $Z$ (or any other single qubit unitary) and building the variational circuit from the gate set $\{R_{X}(\theta),e^{i\phi\left(I\otimes I\otimes Z+I\otimes X\otimes X\right)},e^{i\phi\left(I\otimes I\otimes Z+X\otimes I\otimes X\right)}\}$.
An example circuit for solving the horizontal/vertical line classification problem with a built in symmetry bias is shown in Figure 2. This circuit has a single trainable parameter and solves the problem with 100% accuracy after a standard training procedure.
@@
III Learning Equivariant Maps
In the previous section, we reviewed the use of symmetry in VQCs to reduce the size of the parameter search space. In this section, we tackle the separate problem of learning an equivariant map between two given representations. We do so by leveraging the symmetric VQC construction developed above. Our aim is to build this equivariant map by forcing the estimate, $h_{\theta}$, produced by a new VQC to respect the same $G$-invariance as $f$. This procedure allows us to learn the desired embedding which can then be used for any classification tasks desired. The results that follow focus on this goal.
Given the representation $V:G\to GL(\mathcal{R})$ describing the symmetry respected by the data space and another representation $W:G\to\mathcal{U}(\mathcal{H})$, we can learn an equivariant map between them, assuming one exists. In the context of the symmetry-assisted VQC outlined above, this equivariant map appears as the data embedding, so that we are learning a data embedding which properly respects the symmetry of the data. We begin by embedding the data with a parameterized unitary $\mathcal{E}_{\theta}(x)$ and follow this with a unitary which is $G$-invariant with respect to $W$. The expected value of the circuit will then be $G$-invariant with respect to $V$ whenever the parameterized unitary is an equivariant map between $V$ and $W$ and the observable is $G$-invariant with respect to the $W$ representation. To train the equivariance, we sample from both the data space and the group and minimize the loss function
$L_{g}(x)=(h_{\theta}(V(g)x)-h_{\theta}(x))^{2}\,.$ (28)
Thus, learning an equivariant map amounts to performing a regression minimizing $L_{g}(x)$ for any pair $(g,x)\in G\times\mathcal{R}$. In fact, it suffices to sample over the generators of $G$ since invariance under the action of the generators implies invariance under the action of any other group element. It should be noted that the value of $L_{g}(x)$ can be interpreted as a measure of how equivariant the embedding $\mathcal{E}_{\theta}$ is; that is, how well it transfers the notion of symmetry from one representation to another. In most cases, $L_{g}(x)$ is not actually zero, and so we are really learning maps which are in a sense almost equivariant. By minimizing this loss, we can achieve an embedding that respects the symmetry of our problem. This learned embedding can now be used to construct a symmetric VQC for the purposes of approximating the target function $f$ as discussed in Section II.
In the experiments that follow, we train to 150 epochs with a batch size of 100 and a learning rate of 0.1, and we sample only from a generating set for each group. We simulate our quantum circuits using PennyLane bergholm2018pennylane and optimize using vanilla gradient descent with the gradient computed by PennyLane’s built-in SPSA spall1998overview method, which shifts all parameters of the VQC simultaneously and then approximates the gradient using these shifts and a finite-difference method. The data, which is shuffled at every epoch, is generated in Python.
Example: $C_{2}$ Symmetry
@@ III Learning Equivariant Maps
Consider again the problem of classifying vertical and horizontal lines in a $2\times 2$ image and recall that the classification of a line is invariant under a flip across the vertical axis. Since this flip is its own inverse operation, the group it generates is the cyclic group of order two. Thus, we have a representation $V$ of $C_{2}$ in the dataspace. On the other hand, the operation $\textnormal{SWAP}_{01}\textnormal{SWAP}_{23}$ generates a representation $W$ of $C_{2}$ in the Hilbert space setting. We can attempt to find an equivariant map between these representations by introducing a parameterized embedding $\mathcal{E}_{\theta}(x)=R_{Y}(\theta_{0}+x_{0})\otimes R_{Y}(\theta_{1}+x_{1})\otimes R_{Y}(\theta_{2}+x_{2})\otimes R_{Y}(\theta_{3}+x_{3})$, where $\theta=(\theta_{0},\theta_{1},\theta_{2},\theta_{3})$ and $x=(x_{0},x_{1},x_{2},x_{3})$.
Note that the architecture of the embedding is a free parameter; we have a choice of what gates to allow in its construction. Here, we choose a simple tensor product of rotation gates to easily showcase our procedure. Let $g$ be the generator of $C_{2}$ and observe that $\mathcal{E}_{\theta}\circ V(g)(x)$ is given by
@@
We see that the periodicity of the rotation operators allows for several choices of the $\theta_{i}$ which produce equivariance. For example, when $\theta_{0}=\theta_{1}+2\pi n$ and $\theta_{2}=\theta_{3}+2\pi m$ for some $m,n\in\mathbb{Z}$, it follows that $\mathcal{E}_{\theta}$ is an equivariant map between the representations. Alternatively, if each of the pairs $\theta_{0},\theta_{1}$ and $\theta_{2},\theta_{3}$ differs by an odd multiple of $\pi$, the resulting signs cancel and we again achieve equivariance.
In our simulation we measure with respect to the observable $X\otimes X\otimes Z\otimes Z$, which commutes with the operation $\textnormal{SWAP}_{01}\textnormal{SWAP}_{23}$ and is therefore $G$-invariant, and find that the model learns the parameters $\theta_{0}\approx 3.07\approx\pi$, $\theta_{1}\approx 6.18\approx 2\pi$, $\theta_{2}\approx 4.71\approx 3\pi/2$, and $\theta_{3}\approx 1.65\approx\pi/2$, which aligns with our expectation. The circuit is shown in Figure 3, with the parameter curves and a plot of the average validation loss shown in Figure 4 and Figure 5 respectively.
Example: $C_{2}\times C_{2}$ Symmetry
Let us add an independent symmetry to the mix. The flip $F_{h}$ about the horizontal axis in the $2\times 2$ line classification problem is independent from the vertical flip $F_{v}$, and we will now endeavor to take both into account. The group generated by these independent flips is a direct product of cyclic groups $C_{2}\times C_{2}$. We will make use of the same embedding as for the $C_{2}$ symmetry: $\mathcal{E}_{\theta}(x)=R_{Y}(\theta_{0}+x_{0})\otimes R_{Y}(\theta_{1}+x_{1})\otimes R_{Y}(\theta_{2}+x_{2})\otimes R_{Y}(\theta_{3}+x_{3})$, but this time we expect the parameters to train differently. Indeed, the representation of $C_{2}\times C_{2}$ acting on our Hilbert space will be given by $W(F_{v})=\textnormal{SWAP}_{01}\textnormal{SWAP}_{23}$ and $W(F_{h})=\textnormal{SWAP}_{02}\textnormal{SWAP}_{13}$, so that the unitary applied after the embedding must commute with these two operators. Of course, any choice of a single qubit gate applied simultaneously to all qubits will do the trick. Thus, we take our $C_{2}\times C_{2}$-invariant unitary to be $U=Y^{\otimes 4}$ and we take our observable to be $Z^{\otimes 4}$.
@@
Figure 7: Parameter curves for $C_{2}\times C_{2}$ symmetry.
Figure 8: Average loss for $C_{2}\times C_{2}$ symmetry.
Using the same reasoning as for the $C_{2}$ case, we expect to learn parameters which are equivalent up to some possible phase differences which cancel overall. This is indeed what we find, as the model learns the parameters $\theta_{0}\approx 3.24\approx\pi$, $\theta_{1}\approx 6.22\approx 2\pi$, $\theta_{2}\approx 4.89\approx 3\pi/2$, and $\theta_{3}\approx 4.97\approx 3\pi/2$. The latter two parameters are roughly equivalent, while the former two parameters are about $\pi/2$ from the latter parameter values in opposite directions, thereby canceling signs. See Figure 6 for the circuit, Figure 7 for the parameter curves, and Figure 8 for a plot of the average validation loss.
Example: $D_{4}$ Symmetry
Consider a $3\times 3$ checkerboard distribution of points in the plane. This distribution demonstrates symmetry with respect to the dihedral group $D_{4}$ of order 8. In Figure 9, we plot several sample points which are separated into classes in alternating squares, forming just such a grid. Observe that a rotation $r$ by 90 degrees leaves this classification invariant, as does a flip $F$ about the diagonal line $y=x$. These symmetries together generate $D_{4}$, and so we take this to be our first representation.
@@
For the second representation, we will take the standard representation on four qubits, which is obtained by first identifying $D_{4}$ with the subgroup of the symmetric group $S_{4}$ generated by the permutations (1 2 3 4)1234(1\ 2\ 3\ 4)( 1 2 3 4 ) and (1 2)(3 4)1234(1\ 2)(3\ 4)( 1 2 ) ( 3 4 ) and then identifying these permutations with the corresponding permutation operators which act on four qubits. Thus, for the second representation we take $W(r)=\textnormal{SWAP}_{23}\textnormal{SWAP}_{12}\textnormal{SWAP}_{01}$ and $W(F)=\textnormal{SWAP}_{01}\textnormal{SWAP}_{23}$. Some operators which commute with these generators and are therefore $D_{4}$-invariant include a tensor product of single qubit operators $U^{\otimes 4}$ as well as the permutation $\textnormal{SWAP}_{02}\textnormal{SWAP}_{13}$. Let us therefore apply these operations with $U=Y$ and measure with respect to $Z^{\otimes 4}$. For the architecture of the embedding, we apply $R_{Z}(\theta_{i+4}+x_{(i+1)\!\!\mod 2})R_{Y}(\theta_{i}+x_{i\!\!\mod 2})$ to the $i$-th qubit. See Figure 10 for the circuit.
@@
The validation loss and parameter curves are shown in Figures 12 and 11, respectively. The resulting embedding after training is nearly equivariant. Indeed, taking the data point $(0.40,0.54)$, its rotation, and its reflection, and running the circuit with the trained parameters produces the nearly identical expected values $-0.05429996$, $-0.05429986$, and $-0.05430002$, respectively.
@@
Figure 13: Average loss for $S_{6}$ symmetry.
@@
Example: $S_{6}$ Symmetry
Consider the symmetric group $S_{6}$ on 6 letters. A problem which possesses this symmetry is the classification of graphs with 6 nodes as being either connected or disconnected. A graph is a collection of vertices $V$ and edges $E\subset V\times V$. When $(v,w)\in E$ if and only if $(w,v)\in E$ holds for every $v,w\in V$, we say that the graph is undirected. Otherwise, the graph is called directed or a digraph. We say that a given graph is connected if it has the property that for any vertices $v,w$, there is a sequence of edges $(v_{i},w_{i})$ with $i=0,1,\ldots,n$ such that $v=v_{0}$, $w_{n}=w$, and $w_{i}=v_{i+1}$ for all $i=0,\ldots,n-1$. Intuitively, the graph is connected if there is a path connecting any two vertices. Notice that the labeling of the vertices is arbitrary. We may choose to relabel the vertices with a permutation $\varphi:V\to V$ so long as vertices which are connected by an edge before the permutation remain connected after the permutation (equivalently, the edges are relabeled in precisely the same way). That is, we require that $\varphi$ is a bijection with the property that $(\varphi(v),\varphi(w))$ is an edge if and only if $(v,w)$ is an edge. Such a map is called a graph isomorphism, and it can be shown that it preserves the property of connectedness.
Since permutations of the vertices are equivalent to graph isomorphisms, we take the collection of such maps as our representation of $S_{6}$ in the data space. In the Hilbert space, we take the standard representation of $S_{6}$ given by permuting tensor products according to the given permutation in $S_{6}$. Since this group is generated by the elements (1 2)12(1\ 2)( 1 2 ) and (1 2 3 4 5 6)123456(1\ 2\ 3\ 4\ 5\ 6)( 1 2 3 4 5 6 ), we see that the representation in Hilbert space is given by mapping (1 2)12(1\ 2)( 1 2 ) to $\textnormal{SWAP}_{01}$ and mapping (1 2 3 4 5 6)123456(1\ 2\ 3\ 4\ 5\ 6)( 1 2 3 4 5 6 ) to $\textnormal{SWAP}_{01}\textnormal{SWAP}_{12}\textnormal{SWAP}_{23}\textnormal{SWAP}_{34}\textnormal{SWAP}_{45}$. Certainly a tensor power $U^{\otimes 6}$ of any single qubit operator $U$ commutes with these generators and is therefore $S_{6}$-invariant. For this reason, we take $Z^{\otimes 6}$ as our observable and apply the operation $U=Y^{\otimes 6}$ after the embedding, which we give the following simple architecture: if $(i,j)$ is an edge in the graph with $i\neq j$, apply a controlled rotation with angle $\theta_{i}$ controlled off of qubit $i$ and targeting qubit $j$, and if $i=j$, then simply apply $R_{Y}(\theta_{i})$ to qubit $i$. The circuit is shown in Figure 15 for a particular graph defined by the notation {0: [3, 0, 5, 4], 1: [1, 5, 3, 0, 2], 2: [5, 3, 2, 4], 3: [5], 4: [4, 2], 5: [2, 5, 1, 3]}, which indicates, for example, that there is an edge between vertex 0 and vertices 3, 0, 5, and 4.
The validation loss and parameter curves after training are shown in Figures 13 and 14, respectively. The parameters converged to the same value, which is what we expect since any relabeling of the vertices (maintaining adjacency) gives rise to an equivalent graph.